# Mind the Gap: Data Rewriting for Stable Off-Policy Supervised Fine-Tuning

### v1

@@ Abstract
Supervised fine-tuning (SFT) of large language models can be viewed as an off-policy learning problem, where expert demonstrations come from a fixed behavior policy while training aims to optimize a target policy. Importance sampling is the standard tool for correcting this distribution mismatch, but large policy gaps lead to high variance and training instability. Existing approaches mitigate this issue using KL penalties or clipping, which passively constrain updates rather than actively reducing the gap. We propose a simple yet effective data rewriting framework that proactively shrinks the policy gap by keeping correct solutions as on-policy data and rewriting incorrect ones with guided re-solving, falling back to expert demonstrations only when needed. This aligns the training distribution with the target policy before optimization, reducing importance sampling variance and stabilizing off-policy fine-tuning. Experiments on five mathematical reasoning benchmarks demonstrate consistent and significant gains over both vanilla SFT and the state-of-the-art Dynamic Fine-Tuning (DFT) approach. The data and code will be released at https://github.com/NKU-HLT/Off-Policy-SFT.
@@ 1 Introduction
We propose a simple yet effective data rewriting framework that proactively reduces the policy gap before optimization begins. For each problem, we first sample multiple responses from the target model. If any response solves the problem correctly, we retain it as on-policy data. Otherwise, we prompt the model with the ground-truth solution as a reference to re-solve the problem, generating digest-and-retell data that better reflects the target policy. If both self-solve and re-solve fail, we fall back to the original expert demonstration. Inspired by the intuition that true understanding emerges when learners re-express solutions in their own words rather than copying them verbatim, our digest-and-retell strategy (see Figure 2) transforms the SFT dataset rather than merely constraining optimization dynamics. This process aligns the training distribution more closely with the target policy, reducing importance sampling variance and stabilizing off-policy fine-tuning, with residual mismatch further mitigated through IS weighting.
@@ 5 Conclusion
We introduced a simple yet effective data rewriting framework for supervised fine-tuning of large language models, formulating SFT as an off-policy learning problem. Our approach proactively reduces the policy gap at the data level before training and further alleviates residual mismatch through importance sampling during optimization. Extensive experiments on multiple mathematical reasoning benchmarks demonstrate consistent performance improvements over both vanilla SFT and state-of-the-art Dynamic Fine-Tuning, with the most significant gains observed on base models. These findings underscore the value of data-centric strategies for stabilizing and enhancing off-policy fine-tuning of large language models.

### v2

@@ Abstract
Supervised fine-tuning (SFT) of large language models can be viewed as an off-policy learning problem, where expert demonstrations come from a fixed behavior policy while training aims to optimize a target policy. Importance sampling is the standard tool for correcting this distribution mismatch, but large policy gaps lead to skewed weights, high variance, and unstable optimization. Existing methods mitigate this issue with KL penalties or clipping, which passively restrict updates rather than actively reducing the gap. We propose a simple yet effective data rewriting framework that proactively shrinks the policy gap before training. For each problem, correct model-generated solutions are kept as on-policy data, while incorrect ones are rewritten through guided re-solving, falling back to expert demonstrations only when needed. This aligns the training distribution with the target policy, reducing variance and improving stability. To handle residual mismatch after rewriting, we additionally apply importance sampling during training, forming a two-stage approach that combines data-level alignment with lightweight optimization-level correction. Experiments on five mathematical reasoning benchmarks show consistent and significant gains over both vanilla SFT and the state-of-the-art Dynamic Fine-Tuning (DFT) approach. Data and code will be released at https://github.com/NKU-HLT/Off-Policy-SFT.
@@ 1 Introduction
We propose a simple yet effective data rewriting framework that proactively reduces the policy gap before optimization begins. For each problem, we first sample multiple responses from the target model. If any response solves the problem correctly, we retain it as on-policy data. Otherwise, we prompt the model with the ground-truth solution as a reference to re-solve the problem, generating digest-and-retell data that better reflects the target policy. If both self-solve and re-solve fail, we fall back to the original expert demonstration. Inspired by the intuition that true understanding emerges when learners re-express solutions in their own words rather than copying them verbatim, our digest-and-retell strategy (see Figure 2) transforms the SFT dataset to align it more closely with the target policy.
After rewriting, we apply a lightweight importance sampling correction during training to handle any residual distribution mismatch. This two-stage design—data-level alignment through rewriting combined with optimization-level adjustment via importance sampling—mitigates variance and stabilizes off-policy fine-tuning, ultimately leading to consistent performance gains across benchmarks.
@@ 5 Conclusion
We presented a simple yet effective data rewriting framework for supervised fine-tuning of large language models, viewing SFT through the lens of off-policy learning. Our method proactively reduces the policy gap at the data level before training and further mitigates residual mismatch via importance sampling during optimization. Extensive experiments on multiple mathematical reasoning benchmarks show consistent and significant improvements over both vanilla SFT and the state-of-the-art Dynamic Fine-Tuning approach, with the largest gains achieved on base models. These results highlight the importance of data-centric approaches for stabilizing off-policy optimization and advancing the reasoning capabilities of large language models.